(y–3)^2 = y^2 –9 ????????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (y–3)^2, which is similar to (a–b)^2identity, where (a–b)^2= a^2+b^2-2ab.
(y–3)^2= y^2+(3)^2–2y×3 = y^2+9–6y ≠y^2–9 = RHS
The correct statement is (y–3)^2= y^2+ 9–6y

Thanks